OpenStudy (anonymous):

How do you solve?: cube sqrt 27

OpenStudy (anonymous):

 this is what the problem looks like.

OpenStudy (anonymous):

@Nnesha do you know how to solve this one?

OpenStudy (godlovesme):

i would use a calculator cuz im not good at cube roots :/

OpenStudy (anonymous):

i just need to know how to do it on paper

OpenStudy (godlovesme):

sorry :( idk

OpenStudy (solomonzelman):

you don't need a calculator

OpenStudy (anonymous):

I know. That is what im trying to figure out, how to solve without one

OpenStudy (solomonzelman):

For example, a "Cube root of 8" \(\large\color{slate}{ \sqrt[\color{rd}{3}]{8} }\). There is some \(\large\color{slate}{ {\rm k} }\) that you can multiply three times itself to get 8. \(\large\color{slate}{ ( }\)such that: \(\large\color{slate}{ {\rm k}\times {\rm k}\times{\rm k}=8 }\) \(\large\color{slate}{ ) }\) In this case, you know that 8 is divisible by 2, and when you divide 8 by 2 you get 4. So lets write this. \(\large\color{slate}{ 8=4\times2 }\) now, lets divide 4 by 2, you know that 4 is same as 2 times 2, so: \(\large\color{slate}{ 8=2\times 2\times2 }\) So, this number \(\large\color{slate}{ {\rm k} }\) in ouyr case is \(\large\color{slate}{ 2 }\). Therefore, \(\large\color{slate}{ \sqrt[\color{rd}{3}]{8} =2 }\)

OpenStudy (anonymous):

WELCOME TO OPENSTUDY!!!!

OpenStudy (anonymous):

Great thanks so much

OpenStudy (solomonzelman):

but in your case you need to find this "k" for 27.

OpenStudy (solomonzelman):

what is 27 divisble by?

OpenStudy (anonymous):

9

OpenStudy (solomonzelman):

9?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\large\color{slate}{ 27=9\times 3 }\). but can you reduce the 9?

OpenStudy (anonymous):

cube sqrt 27=9 correct?

OpenStudy (solomonzelman):

9 is a product of which 2 numbers?